# Shaving Weights with Occam’s Razor: Bayesian Sparsification for Neural Networks using the Marginal Likelihood

### v1

@@ Abstract
Neural network sparsification is a promising avenue to save computational time and memory costs, especially in an age where many successful AI models are becoming too large to naïvely deploy on consumer hardware. While much work has focused on different weight pruning criteria, the overall sparsifiability of the network, i.e., its capacity to be pruned without quality loss, has often been overlooked. We present Sparsifiability via the Marginal likelihood (SpaM), a pruning framework that highlights the effectiveness of using the Bayesian marginal likelihood in conjunction with sparsity-inducing priors for making neural networks more sparsifiable. Our approach implements an automatic Occam’s razor that selects the most sparsifiable model that still explains the data well, both for structured and unstructured sparsification. In addition, we demonstrate that the pre-computed posterior Hessian approximation used in the Laplace approximation can be re-used to define a cheap pruning criterion, which outperforms many existing (more expensive) approaches. We demonstrate the effectiveness of our framework, especially at high sparsity levels, across a range of different neural network architectures and datasets.
@@ 1 Introduction
Figure 1: Overview of our proposed SpaM method. We start by training the network to maximize the marginal likelihood using the Laplace approximation, while simplifying the Hessian computation through either the KFAC or diagonal approximation. We can then use our precomputed posterior Hessian as a pruning criterion (OPD). For the unstructured case, we compute thresholds to achieve different target sparsities, compute the mask, and apply it, while for the structured approach, we aggregate the score per layer for easier weight transfer, compute the mask, and then delete the masked structures to obtain a smaller model.
@@
The availability of large datasets and powerful computing infrastructure has fueled the growth of deep learning, enabling the training of massive and complex neural networks. While achieving breakthroughs like high image recognition accuracy (Russakovsky et al., 2015), high-quality text generation (Radford et al., 2022), and catalyzing performance gains across various domains, this development has amplified the challenge of over-parameterization (LeCun et al., 1990; Frankle and Carbin, 2019) and raised concerns about the increase in model size and number of operations. Despite their high performance, over-parameterized neural networks (NNs) present significant deployment challenges, particularly in hardware-constrained environments (Ray, 2022; Paleyes et al., 2022). The quest for sparser neural networks, while promising for efficiency and interpretability, faces a crucial hurdle: many trained networks are not sparsifiable, i.e., they resist effective pruning, regardless of the chosen pruning criteria (Molchanov et al., 2016; Frankle and Carbin, 2019; Liebenwein et al., 2021). Our work tackles this problem, showing that more sparsifiable networks can be achieved through Bayesian model selection using the marginal likelihood (MacKay, 1995; Immer et al., 2021) and the choice of an adequate prior that will induce such sparsifiability.
@@
By leveraging the marginal likelihood in conjunction with well-chosen priors, we use the automatic Occam’s razor property (Rasmussen and Ghahramani, 2000) of Bayesian model selection, guiding the training process towards models that are inherently more sparsifiable while still faithfully representing the data. This is achieved by optimizing thousands of prior parameters to adaptively regularize weight magnitudes. We make use of recent advancements in Laplace inference for Bayesian neural networks (BNNs) (Immer et al., 2021; Daxberger et al., 2021), allowing us to approximate the marginal likelihood (MacKay, 1995) efficiently. Once trained, these BNNs can then be more effectively sparsified using different pruning criteria.
@@
Notably, the pre-computed posterior Hessian from the marginal likelihood training readily translates into a powerful pruning criterion similar to the popular Optimal Brain Damage (OBD; LeCun et al., 1990), which we call Optimal Posterior Damage (OPD), that can be cheaply computed in practice and often outperforms existing criteria.
@@
Extensive empirical evaluations demonstrate the strength of our SpaM approach and the derived OPD pruning criterion in both unstructured and structured sparsification tasks across various datasets, architectures, and sparsification scheduling scenarios (online and post-hoc). Moreover, they show that our framework strikes a compelling balance between performance and computational cost.
@@
  • We propose SpaM, a novel approach to improve the sparsifiability of neural networks using Bayesian model selection with the Laplace-approximated marginal likelihood, and show that it works well in structured and unstructured pruning scenarios, during and after training.
@@
  • We derive OPD, a novel pruning criterion similar to the popular OBD, which outperforms many other criteria in practice and conveniently reuses computations performed in the context of SpaM.
@@ 2 Background
We use deep neural networks to model learning tasks with inputs ${\mathbf{x}}_{n}\in\mathbb{R}^{D}$ and targets ${\mathbf{y}}_{n}\in\mathbb{R}^{C}$ collected in a dataset ${{\mathcal{D}}}=\{({\mathbf{x}}_{n},{\mathbf{y}}_{n})\}_{n=1}^{N}$ of $N$ pairs. A model is parameterized by weights ${\boldsymbol{\theta}}\in\mathbb{R}^{P}$, and maps from inputs to targets using the neural network function ${\mathbf{f}}_{{\boldsymbol{\theta}}}({\mathbf{x}})$. Assuming the data are i.i.d., we have a likelihood $p({{\mathcal{D}}}\mbox{$|$}{\boldsymbol{\theta}})=\prod_{n=1}^{N}p({\mathbf{y}}_{n}\mbox{$|$}{\mathbf{f}}_{{\boldsymbol{\theta}}}({\mathbf{x}}_{n}))$. We minimize the negative log likelihood, which corresponds to common losses like the cross-entropy in classification. Additionally, regularization in the form of weight decay is commonly used and corresponds to a Gaussian prior on parameters $p({\boldsymbol{\theta}})=\mathcal{N}({\boldsymbol{\theta}};{\mathbf{0}},\operatorname{diag}({\boldsymbol{\delta}}))$ with diagonal precision.
@@ 2.1 Marginal Likelihood for Deep Learning
The marginal likelihood serves as the probabilistic foundation for model evaluation and selection. It provides an objective to optimize the tradeoff between data fit and model complexity, akin to the concept of Occam’s razor (Rasmussen and Ghahramani, 2000; MacKay, 2002), by quantifying how well a model ${\mathcal{M}}$, with all its inherent uncertainties, explains the observed data:
@@
The Laplace approximation (LA, MacKay, 1992) provides a tractable and effective approximation to the marginal likelihood for deep learning (Immer et al., 2021). It arises from a second-order Taylor approximation around an estimate of the mode, ${\boldsymbol{\theta}}_{\ast}$, resulting in
$\log p({{\mathcal{D}}}\mbox{$|$}{\mathcal{M}})\approx\log p({{\mathcal{D}}},{\boldsymbol{\theta}}_{\ast}\mbox{$|$}{\mathcal{M}})-\tfrac{1}{2}\log|\tfrac{1}{2\pi}{\mathbf{P}}_{{\boldsymbol{\theta}}_{\ast}}|,$ (2)
@@
In practice, the Hessian of the negative log likelihood is often approximated by the positive semidefinite generalized Gauss-Newton (GGN, Schraudolph, 2002),
${\mathbf{H}}_{{\boldsymbol{\theta}}}\approx{\textstyle\sum_{n=1}^{N}}\nabla_{{\boldsymbol{\theta}}}{\mathbf{f}}_{{\boldsymbol{\theta}}}({\mathbf{x}}_{n})\nabla^{2}_{{\mathbf{f}}}\log p({\mathbf{y}}_{n}\mbox{$|$}{\mathbf{f}}_{{\boldsymbol{\theta}}}({\mathbf{x}}_{n}))\nabla^{\mathrm{\textsf{T}}}_{{\boldsymbol{\theta}}}{\mathbf{f}}_{{\boldsymbol{\theta}}}({\mathbf{x}}_{n}),$ (3)
which relies on the Jacobians of the neural network function and second derivative of the negative log likelihood at the output. Further, it is amenable to efficient structured approximations like diagonal or layer-wise variants (e.g., Martens and Grosse, 2015; Botev et al., 2017).
Diagonal and block-diagonal GGN approximations are efficient and therefore commonly used for Laplace approximations in deep learning (Ritter et al., 2018; Daxberger et al., 2021). The diagonal LA is cheap in terms of storage and computation by only modeling the marginal variances of parameters. Kronecker-factored LA (KFAC LA, Ritter et al., 2018) instead relies on a block-diagonal approximation to the GGN of the parameters ${\boldsymbol{\theta}}_{l}$ in the $l$th layer,
@@
where the factors are given by the outer products of pre-activations and Jacobians w.r.t. the output of a layer, respectively (Martens and Grosse, 2015; Botev et al., 2017). The top left of Figure 1 shows a comparison of both structures.
@@ 2.2 Neural Network Pruning
Unstructured and structured pruning. The goal of the pruning procedure is to remove parameters from ${\boldsymbol{\theta}}$ without affecting the quality of the model output ${\mathbf{f}}_{{\boldsymbol{\theta}}}({\mathbf{x}})$. While unstructured pruning consists in zeroing individual entries $\theta_{p}$ of the weight matrices, structured pruning consists in deleting entire structured sets of parameters $g$, like rows or columns (Liang et al., 2021; Fang et al., 2023). The results of structured pruning enable smaller matrix multiplications which directly provide real-world efficiency gains on most hardware, including GPUs.
Pruning procedures usually follow three steps: (1) We use a scoring function $S(\cdot)$ to evaluate the importance of each individual parameter $S(\theta_{p})$ for unstructured pruning, or structured set of parameters $S(g)$ for structured pruning. (2) We compute a binary mask ${\mathbf{m}}$ with the same dimensions as ${\boldsymbol{\theta}}$ which assign 00 values to parameters whose unstructured or structured pruning scores are below a threshold $T$, and $1$ otherwise. While the threshold $T$ is determined based on the target sparsity across layers for global pruning, it is determined per layer for uniform pruning (Liang et al., 2021). (3) We apply the mask on the weight matrix with element-wise multiplication ${\mathbf{m}}\circ{\boldsymbol{\theta}}$ to effectively remove the least important parameters. Alternatively, structured pruning enables to directly remove rows or columns whose mask values are 00 to reduce weight matrix dimensions.
Pruning before, during, after training. The pruning procedure can be applied at different times. On the one hand, some methods prune before or during training (Lee et al., 2018; Rachwan et al., 2022), thus allowing to (partially) train sparse models. On the other hand, other methods propose to prune after training (LeCun et al., 1990), which allows the compression of existing pre-trained models.
One-shot and iterative pruning. While the pruning procedure can be applied a single time to reach the target sparsity in one-shot manner, it is also possible to apply it iteratively to reach the target sparsity with smaller steps. Further, the pruning iterations can also be spaced at a specified pruning interval during training to gradually compress the model.
@@ 3 Shaving Weights with Occam’s Razor
We identify sparsifiable neural networks by automatically regularizing (groups of) parameters to have small magnitudes to facilitate pruning the least important ones, both within a probabilistic framework. Specifically, we utilize priors that regularize parameters in potentially structured ways leading to smaller magnitudes. To optimize the resulting plentitude of regularization hyperparameters, we employ the Bayesian marginal likelihood as a differentiable objective function, effectively implementing a Bayesian variant of Occam’s razor that drives irrelevant parameters towards smaller magnitudes. While the regularized networks can be pruned with any method, we ultimately propose to use the computed posterior precision for sparsification to only keep well-determined weights of large magnitude.
@@ 3.1 Structured Priors for Regularization
To reduce the magnitude of parameters and make them more amenable to pruning, we introduce structured priors and show how to combine them with diagonal and KFAC Laplace approximations. While a scalar prior, corresponding to weight decay, is the most common, it suggests that all parameters in a neural network are equally relevant and favors uniform magnitude of parameters, which is suboptimal for pruning (Hoefler et al., 2021, Sec. 3.6).
Instead of scalar priors, we regularize parameters with different strengths using layer-, unit-, and parameter-wise priors. Layer-wise priors regularize individual layers differently and have been shown to aid pruning and improve generalization (Aghasi et al., 2017; Gordon et al., 2018; Immer et al., 2021; Antorán et al., 2022). Unit-wise regularization has been used mostly in traditional statistics, for example for group sparsity (Yuan and Lin, 2006), but recently also for channels or feature dimensions in neural networks (Wen et al., 2016; Scardapane et al., 2017).
@@
Our different priors are simple to combine additively with a diagonal Hessian approximation for the Laplace approximation (Equation 2) but not with a KFAC structure. For that reason, so far, only scalar or layer-wise priors have been used for KFAC posterior approximations (Daxberger et al., 2021). The main issue is that we want to preserve the Kronecker factors and not factor them due to the resulting memory cost. For scalar or layer-wise priors, this can be achieved by an eigendecomposition of the individual factors
${\mathbf{A}}\otimes{\mathbf{G}}+{\mathbf{I}}\delta\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}{\mathbf{Q}}_{A}{\boldsymbol{\Lambda}}_{A}{\mathbf{Q}}^{\mathrm{\textsf{T}}}\otimes{\mathbf{Q}}_{G}{\boldsymbol{\Lambda}}_{G}{\mathbf{Q}}_{G}^{\mathrm{\textsf{T}}}+{\mathbf{I}}\delta=({\mathbf{Q}}_{A}\otimes{\mathbf{Q}}_{G})({\boldsymbol{\Lambda}}_{A}\otimes{\boldsymbol{\Lambda}}_{G}+\delta)({\mathbf{Q}}_{A}^{\mathrm{\textsf{T}}}\otimes{\mathbf{Q}}_{G}^{\mathrm{\textsf{T}}}),$ (5)
@@
To add a diagonal prior precision ${\boldsymbol{\delta}}_{l}$ with the KFAC of the $l$th layer, we derive an optimal approximation in the KFAC eigenbasis, so as to maintain the Kronecker-factored structure of the posterior: {restatable}[Diagonal Prior in KFAC Eigenbasis]propositionkfacprop Considering the Frobenius norm, the optimal diagonal perturbation of the KFAC eigenvalues ${\boldsymbol{\Lambda}}_{A}\otimes{\boldsymbol{\Lambda}}_{B}$ to add a diagonal prior precision is given by ${\boldsymbol{\Lambda}}_{A}\otimes{\boldsymbol{\Lambda}}_{B}+\hat{{\boldsymbol{\delta}}}$ with $\operatorname{mat}(\hat{{\boldsymbol{\delta}}})=({\mathbf{Q}}_{G}^{\mathrm{\textsf{T}}})^{2}\mathrm{mat}({\boldsymbol{\delta}}){\mathbf{Q}}_{A}^{2}$ where the square is element-wise and $\mathrm{mat}(\cdot)$ reshapes the vector to match the parameter shape used in KFAC. Thus, it can be computed efficiently without computing a Kronecker product. This approach is similar to that of George et al. (2018), who correct KFAC’s eigenvalues towards the diagonal Gauss-Newton, but solves the problem of adding a full-rank diagonal instead of a rank-1 outer product to the KFAC eigenbasis.

### v2

@@ Abstract
Neural network sparsification is a promising avenue to save computational time and memory costs, especially in an age where many successful AI models are becoming too large to naïvely deploy on consumer hardware. While much work has focused on different weight pruning criteria, the overall sparsifiability of the network, i.e., its capacity to be pruned without quality loss, has often been overlooked. We present Sparsifiability via the Marginal likelihood (SpaM), a pruning framework that highlights the effectiveness of using the Bayesian marginal likelihood in conjunction with sparsity-inducing priors for making neural networks more sparsifiable. Our approach implements an automatic Occam’s razor that selects the most sparsifiable model that still explains the data well, both for structured and unstructured sparsification. In addition, we demonstrate that the pre-computed posterior precision from the Laplace approximation can be re-used to define a cheap pruning criterion, which outperforms many existing (more expensive) approaches. We demonstrate the effectiveness of our framework, especially at high sparsity levels, across a range of different neural network architectures and datasets.
Correspondence to rayen.dhahri@tum.de
@@ 1 Introduction
Figure 1: Overview of our proposed SpaM method. We start by training the network to maximize the marginal likelihood using the Laplace approximation, while simplifying the Hessian computation through either the KFAC or a diagonal approximation. We can then use our precomputed posterior precision as a pruning criterion (OPD). For the case of unstructured pruning, we compute thresholds to achieve different target sparsities, compute the mask, and apply it, while for the structured approach, we aggregate the score per layer for easier weight transfer, compute the mask, and then delete the masked structures to obtain a smaller model.
@@
The availability of large datasets and powerful computing infrastructure has fueled the growth of deep learning, enabling the training of increasingly complex neural networks (NNs). While catalyzing performance gains across various domains, such as image recognition [1] and text generation [2], this development has amplified the challenge of over-parameterization [3, 4] and raised concerns about the increase in model size and computational cost. Over-parameterized neural networks present significant deployment challenges, particularly in hardware-constrained environments [5, 6]. This has sparked the research field of NN sparsification or pruning, where the goal is to remove a (potentially large) number of parameters from a trained network to make it smaller and ultimately cheaper to apply [7, 8]. However, most existing research in this domain has focused on the question of finding better pruning criteria, that is, scoring functions that decide which parameters to prune away [9, 4, 10]. This ignores the challenge that many trained networks are not inherently sparsifiable, i.e., they resist effective pruning, regardless of the chosen criterion. Indeed, standard training of NNs does not encourage sparsifiability at all, so it should not be surprising that such trained NNs would use all of their parameters to some degree to fit the data.
@@
Our work tackles this problem by modifying the training process itself, showing that more sparsifiable networks can be achieved through Bayesian model selection using the marginal likelihood [11, 12] in conjunction with an adequate prior that will induce such sparsifiability. We call this Sparsifiability via the Marginal likelihood, or SpaM. Our approach implements an automatic Occam’s razor [13], guiding the training process towards models that are faithfully fitting the data using only a small subset of their available parameters, such that the remaining ones can be pruned afterwards. This is achieved by optimizing thousands of prior hyper-parameters to adaptively regularize weight magnitudes. We make use of recent advances in Laplace inference for Bayesian neural networks (BNNs) [12, 14], allowing us to approximate the marginal likelihood [11] efficiently.
@@
Once trained, we can use any pruning criterion to more effectively sparsify these networks. Notably, the pre-computed posterior precision of the Laplace approximation obtained from the marginal likelihood training readily translates into a powerful pruning criterion, which we call Optimal Posterior Damage (OPD), similar to the popular Optimal Brain Damage [OBD; 3]. Since it reuses existing computations, it is cheaper than many existing criteria in practice, and it often performs on par or even better.
@@
Extensive empirical evaluations demonstrate the strength of our SpaM approach and the derived OPD pruning criterion in both unstructured and structured sparsification tasks across various datasets and architectures. Moreover, they show that our framework strikes a compelling balance between performance and computational cost.
@@
  • We propose SpaM, a novel approach to improve the sparsifiability of neural networks during training, using Bayesian model selection with the Laplace-approximated marginal likelihood, which works well in structured and unstructured pruning scenarios and with different pruning criteria.
@@
  • We present OPD, a cheap pruning criterion similar to the popular OBD, which performs comparably or better than many other criteria in practice and conveniently reuses computations performed in the context of SpaM.
@@ 2 Background
We use deep neural networks to model learning tasks with inputs ${\mathbf{x}}_{n}\in\mathbb{R}^{D}$ and targets ${\mathbf{y}}_{n}\in\mathbb{R}^{C}$ collected in a dataset ${{\mathcal{D}}}=\{({\mathbf{x}}_{n},{\mathbf{y}}_{n})\}_{n=1}^{N}$ of $N$ pairs. A model is parameterized by weights ${\boldsymbol{\theta}}\in\mathbb{R}^{P}$, and maps from inputs to targets using the neural network function ${\mathbf{f}}_{{\boldsymbol{\theta}}}({\mathbf{x}})$. Assuming the data are i.i.d., we have a likelihood $p({{\mathcal{D}}}\mbox{$|$}{\boldsymbol{\theta}})=\prod_{n=1}^{N}p({\mathbf{y}}_{n}\mbox{$|$}{\mathbf{f}}_{{\boldsymbol{\theta}}}({\mathbf{x}}_{n}))$. We minimize the negative log likelihood, which corresponds to common losses like the cross-entropy in classification. Additionally, regularization in the form of weight decay is commonly used and corresponds to a Gaussian prior on parameters $p({\boldsymbol{\theta}})=\mathcal{N}({\boldsymbol{\theta}};{\mathbf{0}},\operatorname{diag}({\boldsymbol{\delta}}))$ with diagonal precision. We employ Gaussian priors due to their analytical tractability and seamless integration with the Laplace approximation, which requires differentiability. This is essential for maintaining computational efficiency in our framework to approximate the marginal likelihood in a practical and scalable manner. Furthermore, Gaussian priors enable automatic relevance determination by allowing each parameter to have its own variance, facilitating the regularization process without introducing significant computational overhead.
@@ 2.1 Marginal Likelihood for Deep Learning
The marginal likelihood serves as the probabilistic foundation for model evaluation and selection. It provides an objective to optimize the tradeoff between data fit and model complexity, akin to the concept of Occam’s razor [13, 15], by quantifying how well a model ${\mathcal{M}}$, with all its inherent uncertainties, explains the observed data:
@@
The Laplace approximation [LA, 16] provides a tractable and effective approximation to the marginal likelihood for deep learning [12]. It arises from a second-order Taylor approximation around an estimate of the mode, ${\boldsymbol{\theta}}_{\ast}$, resulting in
$\log p({{\mathcal{D}}}\mbox{$|$}{\mathcal{M}})\approx\log p({{\mathcal{D}}},{\boldsymbol{\theta}}_{\ast}\mbox{$|$}{\mathcal{M}})-\tfrac{1}{2}\log|\tfrac{1}{2\pi}{\mathbf{P}}_{{\boldsymbol{\theta}}_{\ast}}({\mathcal{M}})|,$ (2)
@@
In practice, the Hessian of the negative log likelihood is often approximated by the positive semidefinite generalized Gauss-Newton [GGN, 17],
${\mathbf{H}}_{{\boldsymbol{\theta}}}\approx{\textstyle\sum_{n=1}^{N}}\nabla_{{\boldsymbol{\theta}}}{\mathbf{f}}_{{\boldsymbol{\theta}}}({\mathbf{x}}_{n})\nabla^{2}_{{\mathbf{f}}}\log p({\mathbf{y}}_{n}\mbox{$|$}{\mathbf{f}}_{{\boldsymbol{\theta}}}({\mathbf{x}}_{n}))\nabla^{\mathrm{\textsf{\tiny T}}}_{{\boldsymbol{\theta}}}{\mathbf{f}}_{{\boldsymbol{\theta}}}({\mathbf{x}}_{n}),$ (3)
which relies on the Jacobians of the neural network function and second derivative of the negative log likelihood at the output. Further, it is amenable to efficient structured approximations like diagonal or layer-wise variants [e.g., 18, 19].
Diagonal and block-diagonal GGN approximations are efficient, and therefore commonly used for Laplace approximations in deep learning [20, 14]. The diagonal LA is cheap in terms of storage and computation by only modeling the marginal variances of parameters. Kronecker-factored LA [KFAC LA, 20] instead relies on a block-diagonal approximation to the GGN of the parameters ${\boldsymbol{\theta}}_{l}$ in the $l$th layer,
@@
where the factors are given by the outer products of pre-activations and Jacobians w.r.t. the output of a layer, respectively [18, 19]. Here, $A_{l}$ and $G_{l}$ are the uncentered covariances of the respective layer inputs and output gradients. The top left of Figure 1 shows a comparison of both structures.
@@ 2.2 Neural Network Pruning
The goal of the pruning procedure is to remove parameters from ${\boldsymbol{\theta}}$ without affecting the quality of the model output ${\mathbf{f}}_{{\boldsymbol{\theta}}}({\mathbf{x}})$. While unstructured pruning consists in zeroing individual entries $\theta_{p}$ of the weight matrices, structured pruning consists in deleting entire structured sets of parameters $g$, like rows or columns [21, 22]. The results of structured pruning enable smaller matrix multiplications that directly provide real-world efficiency gains on most hardware, including GPUs.
Pruning procedures usually follow three steps: (1) We use a scoring function $S(\cdot)$ to evaluate the importance of each individual parameter $S(\theta_{p})$ for unstructured pruning, or of a structured set of parameters $S(g)$ for structured pruning. (2) We compute a binary mask ${\mathbf{m}}$ with the same dimensions as ${\boldsymbol{\theta}}$, which assigns 00 values to parameters whose unstructured or structured pruning scores are below a threshold $T$, and $1$ otherwise. While the threshold $T$ is determined based on the target sparsity across layers for global pruning, it is determined per layer for uniform pruning [21]. (3) We apply the mask on the weight matrix with element-wise multiplication, ${\mathbf{m}}\circ{\boldsymbol{\theta}}$, to effectively remove the least important parameters. Alternatively, structured pruning enables us to directly remove rows or columns whose mask values are 00 to reduce weight matrix dimensions.
@@ 3 Shaving Weights with Occam’s Razor
We identify sparsifiable neural networks by automatically regularizing (groups of) parameters to have small magnitudes, to facilitate pruning the least important ones, within a probabilistic framework. Specifically, we utilize priors that regularize parameters in potentially structured ways, leading to smaller magnitudes. To optimize the resulting prior hyperparameters, we employ the Bayesian marginal likelihood as a differentiable objective function, effectively implementing a Bayesian variant of Occam’s razor that drives irrelevant parameters towards smaller absolute magnitudes. The regularized networks can then be pruned with any method. However, we additionally propose to reuse the computed posterior precision for sparsification as a cheap and effective criterion.
@@ 3.1 Structured Priors for Regularization
To reduce the magnitude of parameters and make them more amenable to pruning, we introduce structured priors and show how to combine them with diagonal and KFAC Laplace approximations. While a scalar prior, corresponding to weight decay, is the most common, it suggests that all parameters in a neural network are equally relevant and favors a uniform magnitude of parameters, which is suboptimal for pruning [23, Sec. 3.6].
Instead of scalar priors, we regularize parameters with different strengths using layer-, unit-, and parameter-wise priors. Layer-wise priors regularize individual layers differently and have been shown to aid pruning and improve generalization [12, 24, 25, 26]. Unit-wise regularization has been used mostly in traditional statistics, for example, for group sparsity [27], but recently also for channels or feature dimensions in neural networks [28, 29].
@@
Our different priors are simple to combine additively with a diagonal Hessian approximation for the Laplace approximation (Equation 2) but not with a KFAC structure. For that reason, so far, only scalar or layer-wise priors have been used for KFAC posterior approximations [14]. The main issue is that we have to preserve the Kronecker factors to keep the resulting memory cost low. For scalar or layer-wise priors, this can be achieved by an eigendecomposition of the individual factors
${\mathbf{A}}\otimes{\mathbf{G}}+{\mathbf{I}}\delta\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}{\mathbf{Q}}_{A}{\boldsymbol{\Lambda}}_{A}{\mathbf{Q}}_{A}^{\mathrm{\textsf{\tiny T}}}\otimes{\mathbf{Q}}_{G}{\boldsymbol{\Lambda}}_{G}{\mathbf{Q}}_{G}^{\mathrm{\textsf{\tiny T}}}+{\mathbf{I}}\delta=({\mathbf{Q}}_{A}\otimes{\mathbf{Q}}_{G})({\boldsymbol{\Lambda}}_{A}\otimes{\boldsymbol{\Lambda}}_{G}+{\mathbf{I}}\delta)({\mathbf{Q}}_{A}^{\mathrm{\textsf{\tiny T}}}\otimes{\mathbf{Q}}_{G}^{\mathrm{\textsf{\tiny T}}}),$ (5)
@@
To add a diagonal prior precision ${\boldsymbol{\delta}}_{l}$ to the KFAC of the $l$th layer, we derive an optimal approximation in the KFAC eigenbasis, so as to maintain the Kronecker-factored structure of the posterior: {restatable}[Diagonal Prior in KFAC Eigenbasis]propositionkfacprop Considering the Frobenius norm, the optimal diagonal perturbation of the KFAC eigenvalues ${\boldsymbol{\Lambda}}_{A}\otimes{\boldsymbol{\Lambda}}_{B}$ to add a diagonal prior precision is given by ${\boldsymbol{\Lambda}}_{A}\otimes{\boldsymbol{\Lambda}}_{B}+\hat{{\boldsymbol{\delta}}}$ with $\operatorname{mat}(\hat{{\boldsymbol{\delta}}})=({\mathbf{Q}}_{G}^{\mathrm{\textsf{\tiny T}}})^{2}\mathrm{mat}({\boldsymbol{\delta}}){\mathbf{Q}}_{A}^{2}$ where the square is element-wise and $\mathrm{mat}(\cdot)$ reshapes the vector to match the parameter shape used in KFAC. Thus, it can be computed efficiently without computing a Kronecker product.
We provide the proof in Appendix A. The approach is similar to that of George et al. [30], who correct KFAC’s eigenvalues towards the diagonal Gauss-Newton, but solves the problem of adding a full-rank diagonal instead of a rank-1 outer product to the KFAC eigenbasis.